20 years ago Grandma was twice as old as Mum. Today, Grandma is twice the age Mum was when I was born. How old am I?
Answer by ankor@dixie-net.com(22740)

let g = Grandma's age
let m = Mom's age
let y = your age
:
Write an equation for each statement
:
20 years ago Grandma was twice as old as Mum.
g - 20 = 2(m-20)
g - 20 = 2m - 40
g = 2m - 40 + 20
g = 2m - 20
:
Today, Grandma is twice the age Mum was when I was born.
g = 2(m-y)
g = 2m - 2y
:
g = g
2m - 20 = 2m - 2y
Subtract 2m from both sides
-20 = -2y
y = -20/-2
y = 10 yrs old is your age